OpenStudy (anonymous):

Choose all answers that are correct. 4. (1 pt) Which expressions, when evaluated with z = 8, are negative in value? Choose all answers that are correct. A. –2z – 14 B. –25 + 4z C. –2z + 15 D. 30 – 3z

OpenStudy (studygurl14):

Evaluate by plugging in 8 for z in the answer choices -2(8) - 14 = ? -25 + 4(8) = ? -2(8) + 15 = ? 30 - 3(8) = ? Which answer choices result in a negative number for an answer?

alones (alones):

A and C are correct

OpenStudy (anonymous):

a and c?

OpenStudy (studygurl14):

a and C are correct

alones (alones):

yup

OpenStudy (anonymous):

thanks

alones (alones):

you're more than welcomed...

OpenStudy (anonymous):

you both are awesome!:)

alones (alones):

thx you tooo!!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What is the value of –2(–6 + 11) • (–4)?

OpenStudy (anonymous):

IS IT 40?

OpenStudy (anonymous):

???

alones (alones):

Wait i'm solving this one ....

alones (alones):

Yup you got it right !!!

OpenStudy (anonymous):

YEY THANKS

alones (alones):

is their anything else i can solve for you??